# A Comprehensive Survey of Artificial Intelligence Techniques for Talent Analytics

### v2

@@ 6.2 Talent Knowledge Management
6.3 Market-oriented Talent Analytics
@@
AI technology has been effectively applied in labor market analytics [464]. However, those approaches mainly focus on the perspective of global market analysis and have not explored how the changing environment of the labor market affects internal talent management or organization management. In fact, combining macro and micro data in talent analytics is a vital research direction [183]. With the recent accumulation of internal and external data, there exists an exceptional opportunity to implement market-oriented talent analytics. For instance, Hang et al. leveraged the job posting data to capture the potential popularity of employees in external markets specific to skills and further achieve more accurate employee turnover prediction based on the market trend [165]. Moreover, the rapid development of AI technologies provides an excellent technical foundation for this direction. We can unitize multi-task learning [471] to jointly learn both macro and micro talent analytics-related tasks. Heterogeneous graph learning [462, 165] can help us effectively model the correlation of macro and micro data.
@@
Besides, we can use AI technologies to identify various types of talents in market. Identifying high-potential employees has always been an important issue in enterprises. Ye et al. [447] proposed a neural network based dynamic social profiling approach for quantitatively identifying high-potential talents. Cheng et al. [84] conducted a quasi field experiment study and found that AI outperforms humans in the identification of high potential talents. In the labor market, accurate recruitment of digital talents is crucial for the digital transformation of enterprises. Harrigan et al. [166] found a specific relationship between enterprise talent investment and digitization by identifying digital skills-related talents within the organizations. By collecting data on workers, Goos et al. [148] found that people in regular jobs had a harder time finding jobs in new factories compared to digital talents. Babina et al. [30] classified different categories of AI skill-related talents based on job postings on recruitment platforms and analyzed the impact of these talents on the development of the enterprise after entering the enterprise. Kim et al. [211] proposed a dynamic co-occurrence method, which dynamically calculates the AI relevance of various types of skills in the labor market, so as to identify AI talents more accurately. It can be seen that AI has a great impact on human resource management, especially in the current situation where enterprises are paying more and more attention to environmental protection and social responsibility, AI has a broad prospect in the identification of green talents and Environmental, Social and Governance (ESG)-related talents. AI and human resource management still needs large-scale research in the future.
6.4 Organizational Culture Management
@@
In our survey, we reviewed recent advancements in AI techniques for talent analytics in HRM from three perspectives, including talent management, organization management, and labor market analysis. The culture of an organization plays a crucial role in sustaining its effectiveness and viability [151]. Generally, the culture mainly contains three aspects: Mission, Vision, and Values (MVVs), which can help employees understand what is encouraged, discouraged, accepted, or rejected within an organization, and facilitate the organization to thrive with the shared purpose. Recent developments reveal that the availability of extensive datasets covering the entire lifecycle of talents and organizations offers opportunities to realize effective culture management. For example, the interconnection between culture and leadership is evident, with exceptional team leaders significantly shaping organizational culture [151]. Some researchers [227] discussed how ML techniques can be used to inform predictive and causal models of leadership effects. Accordingly, they further provided a step-by-step guide on designing studies that combine field experiments to establish causal relationships with maximal predictive power. Meanwhile, several studies analyze the leadership styles with the data mining algorithms, demonstrating that the different leadership styles significantly influence leadership outcomes [5]. Moreover, there are some researchers who tried to utilize text mining to analyze the organizational culture. For instance, Schmiedel et al. leveraged the online company reviews data and topic model to explore the employee’s perception of corporate culture [357]. Li et al. [235] applied a topic model to obtain firm-level measures of exposure and response related to COVID-19 for many U.S. firms. In detail, they deployed Correlated Topic Model (CTM) [343], which is similar to Latent Dirichlet Allocation (LDA), with 35 topics to discover the correlation between COVID-19 and the company-level measure. The results show that despite the large negative impact of COVID-19 on their operations, firms characterized by strong corporate cultures were able to outperform their counterparts lacking this attribute. As a famous saying goes, “Culture eats strategy for breakfast”, employing AI technologies in organizational culture management will become one of the most critical research directions in the future, as it can help managers scientifically address cultural management.
6.5 Ethical AI in Talent Analytics
@@ 6.5.1 Fairness
The importance of fairness in talent analytics cannot be understated, given its profound impact on employee wellbeing and alignment with organizational values. Although AI technologies have achieved various successes in talent analytics, there is growing concern that such approaches may bring issues of unfairness to people and organizations, as evidenced by some recent reports [228, 101, 50]. For instance, Amazon scraps its AI-based recruitment system due to its discriminatory outcomes against women [101].
@@
Recent studies have delved into the fairness of AI technologies in talent management from different perspectives. For instance, Qin et al. [335] verified that when involving the sensitive features, such as gender, age, etc., into the person-job fit model, the model without special design will easily learn the bias from the original data. Intuitively, we can solve this problem by removing sensitive features, also regarded as one of the pre-processing methods for imposing fairness [98]. However, a large amount of unstructured data already contains sensitive features, such as audio and video data in the interview process. The model can easily infer the potentially sensitive attributes of data, which may still cause the bias of AI algorithms [104]. In their study, Pena et al. examined multimodal systems to predict recruitable candidates using both image and structured data from resumes [323]. The authors first demonstrated that the deep learning model could reproduce the biases from the training data, even without the sensitive features. To address this, they integrated an adversarial regularizer to remove sensitive information from unstructured data and promote algorithmic fairness [295]. Similarly, Yan et al. both leveraged data balancing and adversarial learning to mitigate bias in the multimodal personality assessment  [436]. Moreover, there exist several open-source tools, such as AIF360 [42], FairML [4], Themis-ML [39], that can facilitate systematic bias checks and embed fairness in the AI algorithms [299]. In addition, AI technologies can also help to reduce human bias in different talent management scenarios. For instance, AI technologies have been applied to detect the potentially problematic words in the job posting that lead to bias or even legal risks and further assist employers in writing inclusive job descriptions [275]. Consequently, ensuring the fairness of AI algorithms is becoming an increasingly significant area of research within the field of talent analytics.
6.5.2 Explainability
@@
Recently, there has been an increasing concern among employees and managers regarding the decisions made by black-box AI algorithms. Questions arise regarding the basis of these decisions, understanding the factors behind algorithmic success or failure, and determining how to rectify any errors that occur. Therefore, the research interests in increasing the transparency of AI-based automated decision-making in talent management are re-emerging [187, 470]. For instance, Qin et al. proposed to leverage the attention mechanisms to explain the matching degree between the content of job posting and resume [334]. Zhang et al. further introduced the hierarchical attention and collaborative attention mechanisms to increase the person-job fit model explainability both at the structured and unstructured information level [473]. Upadhyay et al. leveraged the knowledge graph and name entity recognition technologies to generate the understandable textual job recommendation explanation [391]. In  [207], Kaya et al. focused on constructing an end-to-end system for explainable automatic job candidate screening from video resumes. The authors extracted the audio, face, and scene features and leveraged the decision trees to both predict whether the candidates will be invited to the interview and explain the decisions by using binarization with a threshold. Liem et al. further handled the job candidate screening problem from an interdisciplinary viewpoint of psychologists and machine learning scientists [243]. Moreover, Juvitayapun et al. utilized the tree-based model to calculate the importance of different features, enhancing the explainability of AI-based turnover prediction [200].
@@
6.6 Generative AI in Talent Analytics
@@
With technological advances, artificial intelligence has gradually become generative AI, and the language in which AI interacts with humans has changed from “machine language” to “natural language” [420]. By employing suitable templates, all tasks can be addressed through generative AI. Specifically, the knowledge emergence facilitated by large-scale parameters enables generative AI to tackle a series of complex tasks [403]. The advent of generative LLMs have sparked widespread discussion and excitement in both academia and industry. First of all, LLMs have a strong generative ability, which can easily understand various contexts and generate appropriate content according to the prompts. The following work is a good illustration of this aspect of the capabilities of LLMs: Zinjad et al. propose a resume generation tool based on the natural language understanding and information extraction capabilities of LLMs, which allows users to generate tailored personalized resumes by providing simple personal information and job information [490]. Ayoobi et al. propose a method for recognizing LLM-generated resumes in online recruitment platforms, which is beneficial in solving the challenge of identifying fake and LLM-created resumes that are difficult to identify [28]. Magron et al. construct a job posting dataset for skill matching that contains more implicit skills, longer sentences, and is closer to data from real job platforms. They conduct skill-matching experiments using LLMs and show that the method performs well in real-world data evaluation [267]. Second, LLMs facilitate the innovation of traditional AI technologies. For example, in the context of recommender systems, LLMs utilize their high-quality representation of textual features and their extensive coverage of external knowledge to achieve high-quality recommendations [420]. Wu et al. reveal the capability of LLMs to mine graph information by proposing a meta-path prompt constructor to help LLMs understand the semantics of behavioral graphs, and the framework helps provide personalized job recommendations for job seekers [419]. Du et al. improve the traditional LLM-based job recommendation method and verified its effectiveness by mining the explicit and implicit features of users in an online recruitment platform, and aligning the unmatched low-quality and high-quality generated resumes via a generative adversarial network [115]. Zheng et al. guide the LLM-based generative job recommendation system based on the Supervised Fine-Tuning strategy, generate suitable job descriptions according to the resumes of job seekers, and provide individuals with a more personalized and comprehensive job-seeking experience [479]. Abu-Rasheed et al. developed a group chat method based on knowledge graphs to enhance the interpretability of students in conversations with chatbots in the course recommendation task [3].
@@
Besides, in recent years, LLM-based agents are utilized to solve various tasks such as software development, social simulation, and policy simulation [157]. Research on intelligent simulation of management-related problems using digital technologies has also been emerging, such as the study of member autonomy in organizations using human-machine systems based on blockchain technology [119]; how humans can use AI assistants to work effectively with large language models [339]; and how autonomous agents can utilize the generic capabilities of the underlying model for reasoning, decision-making, and environment interaction [441]. Further, by integrating multiple LLM agents, model efficiency can be further improved and simulation effects can be optimized [258]. However, there is not much research on using agents based on large language models for simulating organizational behavior. We should use management theory to simulate the framework of an organization and the behavior of individuals within the organization, which can focus on the trust behavior of inter-organizational personnel in interpersonal interactions, i.e., the willingness to put one’s self-interests at risk based on the positive expectations of others [423]. LLM-based agent can be used to generate conversational data for studying the behaviors and capabilities of chat agents [232]. They query APIs that read and write to web pages, generate content that shapes human behavior, and run system commands as autonomous agents [314]. We can also analyze LLMs’ network formation behavior to examine whether the dynamics of multiple LLMs are similar to or different from human social dynamics [315]. In addition, LLM capabilities can be used to create professional agents with controlled, specialized, and interactive, professional-level capabilities to reshape professional services through evolving expertise [88]. There is an empirical result that shows the agent named QuantAgent is useful in financial research, as it has the capability to uncover viable financial signals and enhance the accuracy of financial forecasts [404]. In a study by Yang et al. [438], an agent-based model was developed to simulate the decision-making behavior and interactive behavior of enterprises and employees based on the HRM characteristics of growing enterprises. The study found that firms should pay extra attention to recruitment programs, changes in workers’ pay gaps are influenced by industry growth and their own capabilities, and pay cap policies have a positive impact on the development of growth firms in the start-up stage. However, there is still a lack of research on Agent-based simulation of cooperative behaviors among employees with different competencies within an organization based on a large language model.
@@
Finally, in the organization, managers use the knowledge inside and outside the organization to make procedural and non-procedural decisions to improve the scientificity and the accuracy of decision making [221]. In order to improve efficiency, many organizational managers have introduced AI to assist in decision making. However, due to the dominant role of humans in decision-making and the lack of autonomy in the interaction of AI, human knowledge and machine knowledge are difficult to function well [468, 212]. The emergence of generative AI gives people hope to solve this problem. Specifically, generative LLMs have the generative properties of being self-adaptive, self-growing, and non-periodic emergence [92, 12]. That is, generative LLMs become more ”proactive” in its interactions with human. In organizational management, generative LLMs can support tasks and decisions of individuals, as well as development planning and work scheduling of the entire organization by organizational managers [57]. For instance, Zheng et al. developed an LLM-based generative job recommendation system to provide individuals with a more personalized and comprehensive job seeking experience [479]. By imposes prompt-based organizational structures on LLM agents, Guo et al. [158] found that LLMs exhibit leadership and spontaneous cooperative behaviors in their organizations. In the decision-making process, it seems that human dominance has become weaker and AI’s decision-making ability has become stronger, which makes it easier to correct the decision-making bias brought about by human intuition [384]. However, due to the algorithmic black box characteristics of generative AI, it is difficult for human to intervene and predict the content generated by generative AI [109]. Therefore, in the decision-making process, generative AI may pose the risk of false information and misleading decisions [368, 408]. Therefore, in talent management, how to dialectically view AI and how to effectively use AI to manage and avoid risks is a key topic worthy of future research. However, the applications of existing LLMs in talent analysis also have limitations, such as the accuracy of generated information, interpretability, interaction forms, etc., which will affect the application of large models in business scenarios
6.7 The Impact of Emergent Events on Talent Analytics

### v3

@@ C.2 Talent Knowledge Management
C.3 Market-oriented Talent Analytics
@@
AI technology has been effectively applied in labor market analytics [229]. However, those approaches mainly focus on the perspective of global market analysis and have not explored how the changing environment of the labor market affects internal talent management or organization management. In fact, combining macro and micro data in talent analytics is a vital research direction [364]. With the recent accumulation of internal and external data, there exists an exceptional opportunity to implement market-oriented talent analytics. For instance, Hang et al. leveraged the job posting data to capture the potential popularity of employees in external markets specific to skills and further achieve more accurate employee turnover prediction based on the market trend [57]. Moreover, the rapid development of AI technologies provides an excellent technical foundation for this direction. We can utilize multi-task learning [365] to jointly learn both macro and micro talent analytics-related tasks. Heterogeneous graph learning [366, 57] can help us effectively model the correlation of macro and micro data.
@@
Besides, we can use AI technologies to identify various types of talents in the market. Identifying high-potential employees has always been an important issue in enterprises. Ye et al. [55] proposed a neural network-based dynamic social profiling approach for quantitatively identifying high-potential talents. Cheng et al. [193] conducted a quasi field experiment study and found that AI outperforms humans in the identification of high-potential talents. In the labor market, accurate recruitment of digital talents is crucial for the digital transformation of enterprises. Harrigan et al. [285] found a specific relationship between enterprise talent investment and digitization by identifying digital skills-related talents within the organizations. By collecting data on workers, Goos et al. [367] found that people in regular jobs had a harder time finding jobs in new factories compared to digital talents. Babina et al. [286] classified different categories of AI skill-related talents based on job postings on recruitment platforms and analyzed the impact of these talents on the development of the enterprise after entering the enterprise. Kim et al. [287] proposed a dynamic co-occurrence method, which dynamically calculates the AI relevance of various types of skills in the labor market, so as to identify AI talents more accurately. It can be seen that AI has a great impact on human resource management, especially in the current situation where enterprises are paying more and more attention to environmental protection and social responsibility, AI has a broad prospect in the identification of green talents and Environmental, Social and Governance (ESG)-related talents. AI and human resource management still need large-scale research in the future.
C.4 Organizational Culture Management
@@
In our survey, we reviewed recent advancements in AI techniques for talent analytics in HRM from three perspectives, including talent management, organization management, and labor market analysis. The culture of an organization plays a crucial role in sustaining its effectiveness and viability [288]. Generally, the culture mainly contains three aspects: Mission, Vision, and Values (MVVs), which can help employees understand what is encouraged, discouraged, accepted, or rejected within an organization, and facilitate the organization to thrive with the shared purpose. Recent developments reveal that the availability of extensive datasets covering the entire lifecycle of talents and organizations offers opportunities to realize effective culture management. For example, the interconnection between culture and leadership is evident, with exceptional team leaders significantly shaping organizational culture [288]. Some researchers [289] discussed how ML techniques can be used to inform predictive and causal models of leadership effects. Accordingly, they further provided a step-by-step guide on designing studies that combine field experiments to establish causal relationships with maximal predictive power. Meanwhile, several studies analyze leadership styles with data mining algorithms, demonstrating that the different leadership styles significantly influence leadership outcomes [290]. Moreover, some researchers try to utilize text mining to analyze the organizational culture. For instance, Schmiedel et al. leveraged the online company reviews data and topic model to explore the employees’ perception of corporate culture [69]. Li et al. [368] applied a topic model to obtain firm-level measures of exposure and response related to emergent events for many U.S. firms. As a famous saying goes, “Culture eats strategy for breakfast”, employing AI technologies in organizational culture management will become one of the most critical research directions in the future, as it can help managers scientifically address cultural management.
C.5 Ethical AI in Talent Analytics
@@ C.5.1 Fairness
The importance of fairness in talent analytics cannot be overstated, given its profound impact on employee well-being and alignment with organizational values. Although AI technologies have achieved various successes in talent analytics, there is growing concern that such approaches may bring issues of unfairness to people and organizations, as evidenced by some recent reports [369, 291, 370]. For instance, Amazon scrapped its AI-based recruitment system due to its discriminatory outcomes against women [291].
@@
Recent studies have delved into the fairness of AI technologies in talent management from different perspectives. For instance, Qin et al. [41] verified that when involving sensitive features, such as gender, age, etc., into the person-job fit model, the model without special design will easily learn the bias from the original data. Intuitively, we can solve this problem by removing sensitive features, also regarded as one of the pre-processing methods for imposing fairness [371]. However, a large amount of unstructured data already contains sensitive features, such as audio and video data in the interview process. The model can easily infer the potentially sensitive attributes of data, which may still cause the bias of AI algorithms [372]. In their study, Pena et al. examined multimodal systems to predict recruitable candidates using both image and structured data from resumes [35]. The authors first demonstrated that the deep learning model could reproduce the biases from the training data, even without the sensitive features. To address this, they integrated an adversarial regularizer to remove sensitive information from unstructured data and promote algorithmic fairness [373]. Similarly, Yan et al. both leveraged data balancing and adversarial learning to mitigate bias in the multimodal personality assessment  [292]. Moreover, there exist several open-source tools, such as AIF360 [374], FairML [375], Themis-ML [376], that can facilitate systematic bias checks and embed fairness in the AI algorithms [377]. In addition, AI technologies can also help to reduce human bias in different talent management scenarios. For instance, AI technologies have been applied to detect the potentially problematic words in the job posting that lead to bias or even legal risks and further assist employers in writing inclusive job descriptions [378]. Consequently, ensuring the fairness of AI algorithms is becoming an increasingly significant area of research within the field of talent analytics.
C.5.2 Explainability
@@
Recently, there has been an increasing concern among employees and managers regarding the decisions made by black-box AI algorithms. Questions arise regarding the basis of these decisions, understanding the factors behind algorithmic success or failure, and determining how to rectify any errors that occur. Therefore, the research interest in increasing the transparency of AI-based automated decision-making in talent management are re-emerging [379, 380]. For instance, Qin et al. proposed to leverage the attention mechanisms to explain the matching degree between the content of job postings and resumes [4]. Zhang et al. further introduced the hierarchical attention and collaborative attention mechanisms to increase the person-job fit model explainability both at the structured and unstructured information level [131]. Upadhyay et al. leveraged the knowledge graph and name entity recognition technologies to generate the understandable textual job recommendation explanation [381]. In  [382], Kaya et al. focused on constructing an end-to-end system for explainable automatic job candidate screening from video resumes. The authors extracted the audio, face, and scene features and leveraged the decision trees to both predict whether the candidates will be invited to the interview and explain the decisions by using binarization with a threshold. Liem et al. further handled the job candidate screening problem from an interdisciplinary viewpoint of psychologists and machine learning scientists [383]. Moreover, Juvitayapun et al. utilized the tree-based model to calculate the importance of different features, enhancing the explainability of AI-based turnover prediction [166].
@@
C.6 Generative AI in Talent Analytics
@@
With technological advances, artificial intelligence has gradually developed a branch called generative AI, and the language in which AI interacts with humans has changed from “machine language” to “natural language” [293]. By employing suitable templates, many tasks can be addressed through generative AI. Specifically, the knowledge emergence facilitated by large-scale parameters enables generative AI to tackle a series of complex tasks [386]. The advent of generative LLMs has sparked widespread discussion and excitement in both academia and industry. First of all, LLMs have a strong generative ability, which enables them to easily understand various contexts and generate appropriate content according to the prompts. The following work is a good illustration of this aspect of the capabilities of LLMs: Zinjad et al. proposed a resume generation tool based on natural language understanding and information extraction capabilities of LLMs, which allows users to generate tailored personalized resumes by providing simple personal information and job information [294]. Ayoobi et al. proposed a method for recognizing LLM-generated resumes in online recruitment platforms, which is beneficial in solving the challenge of identifying fake and LLM-created resumes that are difficult to identify [387]. Magron et al. constructed a job posting dataset for skill matching that contains more implicit skills, longer sentences, and is closer to data from real job platforms. They conduct skill-matching experiments using LLMs and show that the method performs well in real-world data evaluation [356]. Second, LLMs facilitate the innovation of traditional AI technologies. For example, in the context of recommender systems, LLMs utilize their high-quality representation of textual features and their extensive coverage of external knowledge to achieve high-quality recommendations [293]. Wu et al. revealed the capability of LLMs to mine graph information by proposing a meta-path prompt constructor to help LLMs understand the semantics of behavioral graphs, and the framework helps provide personalized job recommendations for job seekers [339]. Du et al. improved the traditional LLM-based job recommendation method and verified its effectiveness by mining the explicit and implicit features of users in an online recruitment platform, and aligning the unmatched low-quality and high-quality generated resumes via a generative adversarial network [136]. Zheng et al. guided the LLM-based generative job recommendation system based on a Supervised Fine-Tuning strategy, generate suitable job descriptions according to the resumes of job seekers, and provide individuals with a more personalized and comprehensive job-seeking experience [135]. Abu-Rasheed et al. developed a group chat method based on knowledge graphs to enhance the interpretability of students in conversations with chatbots in the course recommendation task [388].
@@
Besides, in recent years, LLM-based agents are utilized to solve various tasks such as software development, social simulation, and policy simulation [295]. Research on intelligent simulation of management-related problems using digital technologies has also been emerging, such as the study of member autonomy in organizations using human-machine systems based on blockchain technology [296]; how humans can use AI assistants to work effectively with large language models [389]; and how autonomous agents can utilize the generic capabilities of the underlying model for reasoning, decision-making, and environment interaction [390]. Further, by integrating multiple LLM agents, model efficiency can be further improved and simulation effects can be optimized [297]. However, there is not much research on using agents based on large language models for simulating organizational behavior. We should use management theory to simulate the framework of an organization and the behavior of individuals within the organization, which can focus on the trust behavior of inter-organizational personnel in interpersonal interactions, i.e., the willingness to put one’s self-interest at risk based on the positive expectations of others [298]. An LLM-based agent can be used to generate conversational data for studying the behaviors and capabilities of chat agents [391]. They query APIs that read and write to web pages, generate content that shapes human behavior, and run system commands as autonomous agents [392]. We can also analyze LLMs’ network formation behavior to examine whether the dynamics of multiple LLMs are similar to or different from human social dynamics [29]. In addition, LLM capabilities can be used to create professional agents with controlled, specialized, and interactive, professional-level capabilities to reshape professional services through evolving expertise [393]. There is an empirical result that shows the agent named QuantAgent is useful in financial research, as it has the capability to uncover viable financial signals and enhance the accuracy of financial forecasts [394]. In a study by Yang et al. [395], an agent-based model was developed to simulate the decision-making behavior and interactive behavior of enterprises and employees based on the HRM characteristics of growing enterprises. The study found that firms should pay extra attention to recruitment programs, changes in workers’ pay gaps are influenced by industry growth and their own capabilities, and pay cap policies have a positive impact on the development of growth firms in the start-up stage. However, there is still a lack of research on Agent-based simulation of cooperative behaviors among employees with different competencies within an organization based on a large language model.
@@
Finally, in the organization, managers use the knowledge inside and outside the organization to make procedural and non-procedural decisions to improve the scientificity and accuracy of decision making [396]. To further enhance decision-making efficiency, many organizational managers have introduced AI technologies for assistance. In particular, the recent rise of generative large language models (LLMs) offers new opportunities. These models demonstrate proactive interaction capabilities and can support not only individual decision-making and task execution but also organizational-level planning and scheduling [299]. For instance, Zheng et al. developed an LLM-based generative job recommendation system to provide individuals with a more personalized and comprehensive job seeking experience [135]. By imposing prompt-based organizational structures on LLM agents, Guo et al. [397] found that LLMs exhibit leadership and spontaneous cooperative behaviors in their organizations. In the decision-making process, it seems that human dominance has become weaker and AI’s decision-making ability has become stronger, which makes it easier to correct the decision-making bias brought about by human intuition [398]. However, due to the algorithmic black box characteristics of generative AI, it is difficult for humans to intervene and predict the content generated by generative AI [399]. Therefore, in the decision-making process, generative AI may pose the risk of false information and misleading decisions [400, 401]. Therefore, in talent management, how to dialectically view AI and how to effectively use AI to manage and avoid risks is a key topic worthy of future research. However, the applications of existing LLMs in talent analysis also have limitations, such as the accuracy of generated information, interpretability, interaction forms, etc., which will affect the application of large models in business scenarios.
C.7 The Impact of Emergent Events on Talent Analytics